male-limited precocious puberty results from an autosomal dominate allele (P) that is expressed only in males. If a heterozygous man and a heterozygous woman have children, what proportion of their offspring would be expected to exhibit precocious puberty?
Answer #1

Consider the genotype of heterozygous man and a heterozygous woman is Pp. Cross between these will have the offspring with the following genotype.

Pp* Pp = PP, Pp, Pp, pp

The chances of inheriting dominate allele (P) IS by their offspring (male or female) 3/4.

The chances of inheriting dominant alelle by the male offspring = 1/2* 3/4 = 3/8.

Thus, 3/8 proportion of offspring would be expected to exhibit precocious puberty.